# SpatialBench: Benchmarking Multimodal Large Language Models for Spatial Cognition

### v1

@@ 1 Introduction
In daily life, human can effortlessly integrate spatial information from their surroundings, with a capability known as spatial cognition. This ability extends beyond mere object recognition, serving as a cognitive bridge between perceptual inputs and higher-level functions such as reasoning and navigation. With the rapid advancements of large language models (LLMs) [brown2020language, vicuna2023, driess2023palm, gilardi2023chatgpt, rohan2023alpaca, touvron2023llama], multimodal large language models (MLLMs) have recently emerged as a major step toward general-purpose visual–linguistic intelligence [bai2023qwen, dai2023instructblip, li2023blip, zhang2023internlm, brooks2023instructpix2pix, black2023training, li2023llavamed, zhu2023minigpt, zhang2023gpt4roi, liu2023llava, liu2023improvedllava, ye2023mplug, He2024malmm, Zhang2024groundhog, Chen2023internvl, Yuan2023osprey, Dong2024dreamllm, Cha2023honeybee, qwq32b]. By jointly aligning visual and textual modalities within a shared semantic space, MLLMs have moved beyond abstract visual representations, integrating linguistic context to interpret scenes in a more structured and human-like manner.
Figure 1: In a parking lot scenario, the vehicle must understand the relationships of surrounding objects, reason about possible events, and plan the optimal route to reach the exit.
@@
Recent advances show that MLLMs have exhibited spatial reasoning abilities [Cai2025spatial, Cheng2024spatial, Chen2024spatial, Han2025video, Li2024top, Yamada2024evaluating, zhu2024llava, kumar2025does, yang2023set, tang2024sparkle, wu2025spatial, li2025llava, liu2025oryx], and several benchmarks have been introduced to quantify these capabilities [azuma2022scanqa, ma2023sqa, yang2025thinking, li2025spatial, yin2025spatial, tong2024cambrian, zhang2025from, li2025view]. However, they remain fragmented and task-oriented, often emphasizing performance on specific vision–language tasks rather than assessing spatial cognition as a structured capability. In addition, most benchmarks rely on synthetic or narrowly defined datasets, lacking the visual diversity and real-world complexity necessary to probe genuine spatial cognition. Consequently, these evaluations provide only a partial view of spatial intelligence, making it difficult to analyze cognitive processes and reveal systematic deficiencies in models.
@@
To overcome these limitations, we propose a cognitively grounded evaluation framework for spatial intelligence in MLLMs. Inspired by the cognitive map theory [tolman1948cognitive, o1979precis, battaglia2013simulation] in neuroscience, we conceptualize spatial cognition as a hierarchical process that evolves from low-level perception to high-level reasoning and decision making. Specifically, our framework decomposes spatial understanding into five progressive levels, including observation (L1), topology and relation (L2), symbolic reasoning (L3), causality (L4), and planning (L5), each corresponding to distinct cognitive functions involved in human intelligence. For the example in Figure 1, consider a scenario where a car leaves a parking lot. The model recognizes relevant entities and their spatial configurations (L1), then it understands topological relations such as lane connectivity and obstructions (L2) and maps visual symbols to semantic meanings and evaluates potential detour options (L3). After that, it infers causal outcomes of possible maneuvers (L4), and finally, it integrates prior reasoning to generate a coherent plan (L5). This hierarchical design provides a structured lens through which to interpret model behavior, enabling ability-oriented rather than task-oriented evaluation of spatial intelligence.
@@
Building upon this framework, we construct a large-scale spatial video dataset named SpatialBench that grounds spatial cognition evaluation in realistic multimodal scenarios. Unlike previous synthetic or narrowly scoped datasets, our collection is captured from diverse indoor and outdoor environments, encompassing both static spatial layouts and dynamic scene evolutions that reflect the multi-level cognitive demands of spatial intelligence. To realize the five-level cognitive hierarchy, we design 15 categories of spatial reasoning tasks, each aligned with a distinct stage of spatial cognition. Each video is paired with carefully designed questions and annotations aligned with these cognitive dimensions, enabling systematic, fine-grained, and cognitively interpretable assessment for MLLMs.
@@ 2.1 Multimodal Large Language Models
Recent progress in Large Language Models (LLMs) [brown2020language, vicuna2023, driess2023palm, gilardi2023chatgpt, rohan2023alpaca, touvron2023llama] has catalyzed the evolution of MLLMs, which integrate visual and linguistic modalities within a unified semantic space. By aligning image representations with textual instructions, MLLMs demonstrate remarkable capability in understanding and generating multimodal content across a wide spectrum of real-world tasks [chen2023sharegpt4v, huang2023diversity, alayrac2022flamingo, chen2022pali, driess2023palm]. In general, an MLLM consists of three core components: a modality encoder [radford2021learning, li2022blip], a language backbone (LLM), and a modality interface that connects the two. Recently, the capability of MLLMs has expanded beyond static images to encompass video understanding [chen2024ll3da, fu2024scene, qi2025gpt4scene]. This advancement has led to the incorporation of video–language alignment during pre-training, allowing MLLMs to jointly model temporal semantics and motion dynamics within a unified multimodal framework [lin2024video].
@@ 2.2 Visual-based Spatial Cognition
Visual-based spatial cognition seeks to endow MLLMs with the ability to perceive and reason about three-dimensional spatial relationships directly from visual inputs [Cheng2024spatial, Chen2024spatial, Han2025video, Li2024top, Yamada2024evaluating, mangalam2023egoschema]. Several benchmarks have emerged to evaluate this capability from different perspectives. Video-MME [fu2025videomme] provides a comprehensive assessment across a range of video-related tasks involving recognition and perception. VLM4D [zhou2025vlm4d] emphasizes dynamic motion analysis, and STI-Bench [li2025sti] examines physical reasoning by testing models’ ability to predict and estimate object motions and displacements. VSI-Bench [yang2025thinking] introduces a structured benchmark comprising eight question types, each designed to probe different dimensions of spatial understanding in MLLMs. Ego-ST Bench [wu2025st] extends this to self-centered navigation in egocentric environments. SpatialLadder [li2025spatial] presents a comprehensive dataset spanning multiple categories, covering spatial reasoning tasks from single-image understanding to video-based inference. In contrast, MindCube [yin2025spatial] assesses MLLMs’ ability to infer complete spatial structures from limited visual observations.
Existing spatial understanding benchmarks lack a unified, ability-oriented framework, focusing instead on isolated tasks such as recognition, grounding, or motion prediction. Their evaluation design remains fragmented, failing to capture the hierarchical progression of spatial cognition from perception to reasoning. Moreover, most datasets are constructed by extending simple indoor scenes [yang2025thinking, wu2025spatial], offering limited diversity in environmental dynamics and spatial complexity. As a result, they provide only partial insights into MLLMs’ spatial reasoning capabilities, leaving the broader question of systematic, cognitively grounded evaluation largely unexplored.
@@ 3.1 Conceptual Foundations of Spatial Cognition
Spatial cognition encompasses the processes that enable intelligent systems to perceive, represent, and reason about spatial relationships within their environment. It involves acquiring spatial knowledge from sensory inputs, forming internal representations of space, and utilizing these representations for high-level tasks [tolman1948cognitive, kozhevnikov2023different]. The cognitive map theory [tolman1948cognitive, o1979precis] provides a foundational opinion of how such representations are organized. Originating from behavioral studies of animals and later supported by discoveries of place cells in the hippocampal system, this theory proposes that intelligent agents construct internal, map-like structures that encode both metric and topological relations. These cognitive maps allow flexible navigation, route planning, and generalization beyond direct sensory experience. More importantly, they reveal that spatial knowledge is not a flat or static representation, but a hierarchically organized system, where low-level perceptual and motor cues are progressively abstracted into higher-order representations that integrate semantic, relational, and causal information. Building on this foundation, modern computational perspectives extend spatial cognition toward causal and multimodal understanding [battaglia2013simulation, lake2017building, botvinick2017building]. This hierarchical and integrative view provides the theoretical basis for our proposed framework, which conceptualizes spatial cognition in MLLMs as a progressive process evolving from perception to reasoning and ultimately to planning.
@@ 3.2 Hierarchical Spatial Cognitive Taxonomy
Based on the cognitive map theory, we propose the first systematic and hierarchical framework for spatial cognition evaluation, capturing the progressive development of spatial understanding from perception to high-level reasoning. Unlike prior benchmarks that focus on isolated visual tasks, our taxonomy is ability-driven, which means that each level represents a distinct, measurable cognitive capacity that reflects a specific stage of spatial intelligence. This hierarchical framework delineates spatial cognition into five progressive levels, each corresponding to a fundamental stage in the transition from sensory perception to deliberative reasoning: observation (L1), topology and relation(L2), symbolic reasoning (L3), causality (L4), and planning (L5). Together, these levels illustrate the progressive process by which intelligent systems transform raw perceptual information into organized spatial reasoning.
@@
Causality. The causality level reflects the ability to infer spatiotemporal dependencies and predict outcomes of actions. It involves reasoning about how object movements, physical interactions, or agent behaviors lead to specific spatial consequences, integrating physical intuition and causal understanding into spatial reasoning.
Planning. This level represents the highest stage of spatial cognition, where perception, relational understanding, and causal reasoning are integrated to enable deliberate, goal-oriented decision making. At this level, an agent should synthesize its spatial representations and predictive reasoning to formulate coherent action sequences or navigation strategies that adapt to dynamic environmental contexts.
Collectively, these five levels delineate a progressive hierarchy that mirrors the cognitive evolution from sensory perception to reasoning and planning. By organizing spatial understanding into an ability-driven, hierarchically structured framework, our taxonomy provides the first systematic and cognitively grounded paradigm for assessing spatial intelligence in MLLMs, offering a unified foundation for interpreting and benchmarking spatial cognition across models of varying architectures and scales.
@@ 4.1 Overview
We introduce SpatialBench, a large-scale benchmark for assessing evaluate the hierarchical spatial cognition of MLLMs using first-person videos. The dataset comprises 15 question types, each carefully mapped to one of the five cognitive levels introduced above:
@@
The SpatialBench dataset comprises 1,347 question–answer pairs sourced from 50 videos captured from an egocentric perspective. The collection spans both indoor and outdoor settings and includes static as well as dynamic scenes, covering challenging real-world contexts such as city roads, forest trails, residential areas, and underground environments. Together, these recordings provide diverse temporal and spatial complexity for evaluating MLLMs’ spatial cognition ability in realistic scenarios. The detailed statistics of SpatialBench are shown in Figure 4.
Figure 4: The statistics of SpatialBench.
@@ 4.2 Benchmark Construction
Annotation Verification. We implement a multi-step verification protocol to ensure the reliability of the generated annotations. For L1 and L2 questions, multiple leading models independently generate answers, and the consistency of these responses is evaluated. Questions with fully consistent model outputs are provisionally approved, but a subset of these automatically approved answers is further subject to human spot-checking to guarantee overall quality. Any discrepancies detected, whether during consistency checks or spot audits, trigger full human review. For question beyond L3, all annotations undergo mandatory human verification due to their higher cognitive complexity. The human review process follows a fixed checklist including: whether the evidence frames display key entities, whether multi-model outputs have been correctly interpreted, and whether answers conform to the predefined question schema. Annotations that are modified or rejected during the first review are subsequently evaluated by an additional annotator. This annotator examines the evidence summaries, the explanations and supporting screenshots submitted by the human reviewers, and also inspects the video and the question directly. Based on this comprehensive review, this annotator determines the final annotation. The verification protocol used ensures that all QA annotations are accurate, consistent, auditable, and reproducible, providing a trustworthy foundation for evaluating MLLMs’ spatial cognition.
Figure 5: Comparison of models across spatial cognitive levels.
@@
Model Ave. Score Rank Observation Topology & Relation Symbolic Reasoning Causality Planning
Random 25.06 25.32 - - - - - 24.95 24.84 25.76 24.62 24.67 25.24 24.63 24.98 25.57 24.60 25.64
- 24.97 24.95 25.57 25.50
@@
Gemini-2.5-pro 71.86 75.79 1 57.13 72.78 48.26 38.66 69.39 62.24 78.79 62.24 91.00 73.20 94.90 86.46 80.81 73.33 74.23
55.63 72.82 84.88 80.81 74.11
GPT-4o-mini 33.44 30.92 9 31.49 54.17 28.26 24.95 73.47 49.48 26.53 34.69 20.00 31.96 9.18 23.96 33.33 0.00 31.96
36.12 40.73 21.65 33.33 27.68
Claude-sonnet-4-5 32.10 28.56 12 41.59 70.51 40.87 41.67 21.79 13.58 32.10 33.33 20.99 36.25 30.00 26.58 22.50 20.00 28.75
50.27 24.38 30.96 22.50 27.78
GPT-5-chat-latest 23.93 22.45 17 38.33 72.15 42.61 36.48 5.32 3.19 29.47 17.02 19.79 16.13 20.43 10.87 20.00 13.33 25.81
48.48 15.01 15.83 20.00 24.07

### v3

@@ 1 Introduction
In daily life, human can effortlessly integrate spatial information from their surroundings, with a capability known as spatial cognition. This ability extends beyond mere object recognition, serving as a cognitive bridge between perceptual inputs and higher-level functions such as reasoning and navigation. With the rapid advancements of large language models (LLMs) [9, 18, 21, 24, 55, 60], multimodal large language models (MLLMs) have recently emerged as a major step toward general-purpose visual–linguistic intelligence [4, 19, 39, 71, 8, 6, 36, 77, 72, 45, 44, 67, 27, 73, 16, 69, 20, 11, 57]. By jointly aligning visual and textual modalities within a shared semantic space, MLLMs have moved beyond abstract visual representations, integrating linguistic context to interpret scenes in a more structured manner. Recent advances show that MLLMs have exhibited spatial reasoning abilities [10, 17, 12, 26, 35, 64, 76, 53, 65, 54, 61, 34, 46], and several benchmarks have been introduced to quantify these capabilities [3, 47, 66, 38, 68, 59, 70, 37]. However, they remain fragmented and task-oriented, often emphasizing performance on specific vision–language tasks rather than assessing spatial cognition as a structured capability. In addition, most benchmarks rely on synthetic or narrowly defined datasets, lacking the visual diversity and real-world complexity necessary to probe genuine spatial cognition. Consequently, these evaluations provide only a partial view of spatial intelligence, making it difficult to analyze cognitive processes and reveal systematic deficiencies in models.
Figure 1: In a parking lot, the vehicle must understand the relationships of surrounding objects, reason about possible events, and plan the optimal route to reach the exit.
@@
To overcome these limitations, we propose a cognitively grounded evaluation framework for spatial intelligence in MLLMs. Inspired by the cognitive map theory [58, 49, 5] in neuroscience, we conceptualize spatial cognition as a hierarchical process that evolves from low-level perception to high-level reasoning and decision making. Specifically, our framework decomposes spatial understanding into five progressive levels, including observation (L1), topology and relation (L2), symbolic reasoning (L3), causality (L4), and planning (L5), each corresponding to distinct cognitive functions involved in human intelligence. For the example in Fig. 1, consider a scenario where a car leaves a parking lot. The model recognizes relevant entities and their spatial configurations (L1), then it understands topological relations such as lane connectivity and obstructions (L2) and maps visual symbols to semantic meanings and evaluates potential detour options (L3). After that, it infers causal outcomes of possible maneuvers (L4), and finally, it integrates prior reasoning to generate a coherent plan (L5). This hierarchical design provides a structured lens through which to interpret model behavior, enabling ability-oriented rather than task-oriented evaluation of spatial intelligence.
@@
Building upon this framework, we construct a large-scale spatial video dataset named SpatialBench that grounds spatial cognition evaluation in realistic scenarios. Unlike previous synthetic or narrowly scoped datasets, our collection is captured from diverse indoor and outdoor environments, encompassing both static spatial layouts and dynamic scene evolutions that reflect the multi-level cognitive demands of spatial intelligence. To realize the five-level cognitive hierarchy, we design 15 categories of spatial reasoning tasks, each aligned with a distinct stage of spatial cognition. Each video is paired with carefully designed questions and annotations aligned with these cognitive dimensions, enabling systematic, fine-grained, and cognitively interpretable assessment for MLLMs.
@@ 2.1 Multimodal Large Language Models
Recent progress in Large Language Models (LLMs) [9, 18, 21, 24, 55, 60] has catalyzed the evolution of MLLMs, which integrate visual and linguistic modalities within a unified semantic space. By aligning image representations with textual instructions, MLLMs demonstrate remarkable capability in understanding and generating multimodal content across a wide spectrum of real-world tasks [13, 29, 2, 15, 21]. In general, an MLLM consists of three core components: a modality encoder [52, 40], a language backbone (LLM), and a modality interface that connects the two. Recently, the capability of MLLMs has expanded beyond static images to encompass video understanding [14, 23, 50]. This advancement has led to the incorporation of video–language alignment during pre-training, allowing MLLMs to jointly model temporal semantics and motion dynamics within a unified multimodal framework [42].
@@ 2.2 Visual-based Spatial Cognition
Visual-based spatial cognition seeks to endow MLLMs with the ability to perceive and reason about three-dimensional spatial relationships directly from visual inputs [17, 12, 26, 35, 64, 48]. Several benchmarks have emerged to evaluate this capability from different perspectives. VSP [63] highlights the bottleneck of perception and reasoning in spatial planning, focusing on static scenes with picture inputs. Video-MME [22] provides a comprehensive assessment across a range of video-related tasks involving recognition and perception. VLM4D [75] emphasizes dynamic motion analysis, and STI-Bench [41] examines physical reasoning by testing models’ ability to predict and estimate object motions and displacements. VSI-Bench [66] introduces a structured benchmark comprising eight question types, each designed to probe different dimensions of spatial understanding in MLLMs. Ego-ST Bench [62] extends this to self-centered navigation in egocentric environments. SpatialLadder [38] presents a comprehensive dataset spanning multiple categories, covering spatial reasoning tasks from single-image understanding to video-based inference. MindCube [68] assesses MLLMs’ ability to infer complete spatial structures from limited visual observations.
Benchmark Data Nature Modality Perspectives Taxonomy Scale (Vid/QA)
VSP Synthetic & 2D images Image - 4 Levels - / 4.4k
VLM4D Hybrid Video Ego & Exo Flat 1k / 1.8k
Video-MME Mixed Video Mixed Flat 900 / 2.7k
VSI-Bench Open Source Datasets (Indoor) Video Ego 3 Levels 288 / 5k+
STI-Bench Real Video Ego & Exo 2 Levels 300 / 2k
SpatialBench (Ours) Self-recorded (In/Outdoor) Video Ego & Exo 5 Levels 117 / 3.1k+
Table 1: Comparison of SpatialBench with existing spatial and video benchmarks.
While simulation-based benchmarks like VSP offer controlled environments, they often fail to capture the physical nuances of the real world. Other real-life video based benchmarks like VSI-Bench utilize orderly and open-source indoor videos, lacking noise interference in practical application scenarios (such as outdoor robots). SpatialBench distinguishes itself by featuring diverse real-world video sequences across both indoor and outdoor settings, challenging MLLMs with complex lighting, sensor noise, and authentic 3D geometry that are absent in synthetic data. We present the differences between SpatialBench and other benchmarks in Table 1.
@@ 3.1 Conceptual Foundations of Spatial Cognition
Spatial cognition encompasses the processes that enable intelligent systems to perceive, represent, and reason about spatial relationships within their environment. It involves acquiring spatial knowledge from sensory inputs, forming internal representations of space, and utilizing these representations for high-level tasks [58, 31]. The cognitive map theory [58, 49] provides a foundational opinion of how such representations are organized. Originating from behavioral studies of animals and later supported by discoveries of place cells in the hippocampal system, this theory proposes that intelligent agents construct internal, map-like structures that encode both metric and topological relations. These cognitive maps allow flexible navigation, route planning, and generalization beyond direct sensory experience. More importantly, they reveal that spatial knowledge is not a flat or static representation, but a hierarchically organized system, where low-level perceptual and motor cues are progressively abstracted into higher-order representations that integrate semantic, relational, and causal information. Building on this foundation, modern computational perspectives extend spatial cognition toward causal and multimodal understanding [5, 32, 7]. This hierarchical and integrative view provides the theoretical basis for our framework, which conceptualizes spatial cognition in MLLMs as a progressive process evolving from perception to reasoning and ultimately to planning.
@@ 3.2 Hierarchical Spatial Cognitive Taxonomy
Based on the cognitive map theory, we propose the first systematic and hierarchical framework for spatial cognition evaluation, capturing the progressive development of spatial understanding from perception to high-level reasoning. Unlike prior benchmarks that focus on isolated visual tasks, our taxonomy is ability-driven, where each level represents a distinct, measurable cognitive capacity that reflects a specific stage of spatial intelligence. This hierarchical framework delineates spatial cognition into five progressive levels, each corresponding to a fundamental stage in the transition from sensory perception to deliberative reasoning: observation (L1), topology and relation(L2), symbolic reasoning (L3), causality (L4), and planning (L5). Together, these levels illustrate the progressive process by which intelligent systems transform raw perceptual information into organized spatial reasoning, as shown in Fig. 2.
@@
Causality. This level reflects the ability to infer spatiotemporal dependencies and predict outcomes of actions. It involves reasoning about how object movements, physical interactions, or agent behaviors lead to specific consequences, integrating physical intuition and causal understanding into reasoning.
Planning. It represents the highest stage of spatial cognition, where perception, relational understanding, and causal reasoning are integrated to enable deliberate, goal-oriented decision making. At this level, an agent should synthesize its spatial representations and predictive reasoning to formulate coherent action sequences or navigation strategies that adapt to dynamic environmental contexts.
Collectively, these five levels delineate a progressive and interdependent hierarchy, where foundational sensory capabilities serve as the indispensable bedrock for higher-order reasoning and planning. This structural dependency marks a significant departure from existing benchmarks, which predominantly evaluate model efficacy through the narrow lens of isolated, fragmented tasks. Rather than yielding disjointed performance metrics, our taxonomy enables researchers to pinpoint the precise developmental stage of a model’s spatial cognition and diagnose its cognitive bottlenecks. By organizing spatial understanding into this ability-driven, cumulative continuum, we provide the first systematic and cognitively grounded paradigm. It offers a unified foundation not merely for benchmarking, but for interpreting the evolutionary trajectory of spatial intelligence across MLLMs of varying architectures and scales.
Figure 3: An overview of SpatialBench.
@@ 4.1 Overview
We introduce SpatialBench, as shown in Fig. 3, a large-scale benchmark for assessing evaluate the hierarchical spatial cognition of MLLMs using first-person videos. The dataset comprises 15 question types, each carefully mapped to one of the five cognitive levels introduced above:
@@
The SpatialBench dataset comprises 3193 question–answer pairs sourced from 117 videos captured by us in real life. The collection spans both indoor and outdoor settings and includes static as well as dynamic scenes, covering challenging real-world contexts such as city roads, forest trails, residential areas, and underground environments. Together, these recordings provide diverse temporal and spatial complexity for evaluating MLLMs’ spatial cognition ability in realistic scenarios. The detailed statistics of SpatialBench are shown in Fig. 4.
@@ 4.2 Benchmark Construction
Annotation Verification. We implement a multi-step verification protocol to ensure the reliability of the generated annotations. For L1 and L2 questions, multiple leading models independently generate answers, and the consistency of these responses is evaluated. Questions with fully consistent model outputs are provisionally approved, but a subset of these automatically approved answers is further subject to human spot-checking to guarantee overall quality. Any discrepancies detected, whether during consistency checks or spot audits, trigger full human review. For question beyond L3, all annotations undergo mandatory human verification due to their higher cognitive complexity. The human review process follows a fixed checklist including: whether the evidence frames display key entities, whether multi-model outputs have been correctly interpreted, and whether answers conform to the predefined question schema. Annotations that are modified or rejected during the first review are subsequently evaluated by an additional annotator. This annotator examines the evidence summaries, the explanations and supporting screenshots submitted by the human reviewers, and also inspects the video and the question directly. Based on this comprehensive review, this annotator determines the final annotation. The verification protocol used ensures that all QA annotations are accurate, consistent, auditable, and reproducible, providing a trustworthy foundation for evaluation.
Model Score (Rank) Ave. Obs. Topo. Symb. Caus. Plan.
Human Level 96.40 (-) 93.89 77.78 100.00 100.00 100.00 100.00
Random 25.32 (-) 25.06 24.95 24.97 24.95 25.57 25.50
Proprietary Models
Gemini-2.5-pro 75.79 (1) 71.86 55.63 72.82 84.88 80.81 74.11
GPT-4o-mini 30.92 (9) 33.44 36.12 40.73 21.65 33.33 27.68
Claude-sonnet-4-5 28.56 (12) 32.10 50.27 24.38 30.96 22.50 27.78
GPT-5-chat-latest 22.45 (17) 23.93 48.48 15.01 15.83 20.00 24.07
Open Source Models
Qwen3-VL-235B-A22B-Inst 37.79 (2) 38.84 48.68 37.40 32.99 33.33 41.07
Qwen3-VL-235B-A22B-Thin 36.27 (3) 36.60 47.70 33.33 29.55 28.28 44.64
Qwen2.5-VL-72B-Inst 33.82 (4) 32.73 38.33 29.01 33.33 36.36 33.04
Qwen3-VL-30B-A3B-Inst 32.83 (7) 37.07 48.52 41.78 21.31 29.29 33.04
Qwen3-VL-30B-A3B-Thin 31.22 (8) 36.83 46.20 41.46 26.12 26.26 28.57
GLM-4V-Plus-0111 30.40 (11) 36.89 49.81 40.90 23.71 36.73 17.86
ERNIE-4.5-Turbo-VL-32K 28.05 (13) 34.72 44.25 41.58 20.62 30.30 18.75
GLM-4.5V 27.72 (14) 35.34 50.16 38.21 25.00 28.12 16.96
MiniMax-Text-01 27.59 (15) 33.44 48.14 35.90 22.34 28.28 19.64
QwQ-32B 23.81 (16) 29.76 38.30 34.89 18.21 28.28 13.39
Qwen2.5-VL-7B-Instruct 17.45 (18) 17.84 19.69 15.48 17.96 18.95 16.22
Table 2: Comprehensive evaluation results on SpatialBench. Models are grouped by category and sorted by their overall Rank. Bold indicates the best performance overall; underline indicates the best performance among open-source models. For presentation, we simplified the table.
@@
Model Observation Topology & Relation Symbolic Reasoning Causality Planning
Human Level 100.00 53.75 67.50 60.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Random - - - - 24.95 24.84 25.76 24.62 24.67 25.24 24.63 24.98 25.57 24.60 25.64
@@
Gemini-2.5-pro 57.13 72.78 48.26 38.66 69.39 62.24 78.79 62.24 91.00 73.20 94.90 86.46 80.81 73.33 74.23
GPT-4o-mini 31.49 54.17 28.26 24.95 73.47 49.48 26.53 34.69 20.00 31.96 9.18 23.96 33.33 0.00 31.96
Claude-sonnet-4.5 41.59 70.51 40.87 41.67 21.79 13.58 32.10 33.33 20.99 36.25 30.00 26.58 22.50 20.00 28.75
GPT-5-chat-latest 38.33 72.15 42.61 36.48 5.32 3.19 29.47 17.02 19.79 16.13 20.43 10.87 20.00 13.33 25.81